### v2

Abstract
Peer influence on effort devoted to some activity is often studied when effort is unobserved and the researcher instead observes an outcome that combines effort with other shocks. For instance, in education, achievement measures such as GPA reflect both effort and idiosyncratic GPA shocks. We propose an alternative approach that circumvents this approximation. Our framework distinguishes unobserved shocks to GPA that do not affect effort from preference shocks that do affect effort levels. We show that peer effects estimates obtained using our approach can differ significantly from classical estimates (where effort is approximated) if the network includes isolated students. Applying our approach to data on high school students in the United States, we find that peer effect estimates relying on GPA as a proxy for effort are 40% lower than those obtained using our approach.
Keywords: Social networks, Peer effects, Academic achievement, Unobserved effort, Isolated agents
JEL classification: C31, J24
1  Introduction
In recent years, there has been a growing interest in understanding how peers impact educational outcomes (sacerdote2011peer; epple2011peer). Because peer effects on students’ academic effort may involve a social multiplier effect, understanding whether students are influenced by their friends and the size of this influence is crucial for evaluating policies aimed at improving academic achievement (m1993). However, estimating peer effects based on academic effort presents challenges, since effort itself is generally unobserved. Consequently, many empirical studies estimate peer effects on achievement (test scores or GPA) as the policy-relevant outcome, often grounding this choice in models where achievement depends on own and peers’ effort. While microfounded models exploring the impact of peer interactions on academic achievement are often focused on effort, GPA is often the main observed achievement measure in the empirical analysis, especially in settings where direct measures of student effort are limited or unavailable (e.g., see cpz2009; arcidiacono2012estimating; fruehwirth2013identifying; fruehwirth2014can; hong2017sitting). Yet, this approximation overlooks that GPA is not solely influenced by effort but also by many other factors, including unobserved student and school characteristics.
In this paper, we investigate the implications of using GPA as a proxy for academic effort on the estimation and interpretation of peer effects. We develop a structural model of educational effort with social interactions, in which students decide on their academic effort while taking into account that of their peers (friends). We also explicitly model the production function of academic achievement (GPA), which includes effort as a key input. Without using a proxy for student academic effort, we show that our model identifies peer effects on academic effort. Furthermore, we demonstrate that the size of the effect may differ from that estimated using GPA as a proxy for effort, because the proxy-based approach suffers from an issue akin to an omitted variables bias. From the structural model, we derive a reduced-form equation for GPA, which differs from the classical linear-in-means peer effects specification, that highlights the need to control for two types of unobserved GPA shocks to disentangle peer effects on academic effort from other common effects captured by GPA. First, one needs to account for common shocks that directly influence GPA, such as improvements in teaching quality, irrespective of the effort level. These shocks result in a GPA increase for the same level of effort and do not involve a social multiplier effect. Second, one needs to account for common shocks affecting students’ preferences, such as increasing motivation to value academic achievement through information, which influence both academic effort and GPA, and may have social multiplier effects among students with friends.
We demonstrate that approximating student effort with GPA may result in biased estimates of peer effects when some students in the network are isolated (i.e., have no friends). This is because students who have friends are affected differently by the two types of GPA shocks mentioned above compared to isolated students. Standard approaches that use GPA as a proxy fail to differentiate between these two types of shocks, leading to biased estimates of peer effects when the network includes isolated students. However, we show that in networks without isolated students, the standard model does not produce biased estimates of peer effects. The key difference between the reduced-form of the standard model and our framework is that while the standard model has a single intercept for each school, our approach accounts for unobserved school-level heterogeneity based on whether a student has friends or not. This amounts to introducing two types of fixed effects per school, one for students with friends and another for isolated students. This two‐fixed‐effect approach is uncommon in linear‐in‐means models—one exception is boucher2024toward, who use it for different identification purposes. [^1] Their analysis requires identifying peer effects from the equation for isolated individuals. As a result, they treat the GPA equation separately for isolated and non-isolated nodes.
Our econometric model incorporates two types of school-level unobserved shocks to GPA, which renders the well-known reflection problem more complex. [^2] The reflection problem arises when one cannot disentangle endogenous peer effects from exogenous contextual peer effects (m1993) In the case of the standard linear-in-means model, BramoulleDjebbariFortin2009 provide straightforward conditions relating to the network structure under which the reflection problem is resolved. We extend their identification analysis to our framework; our main condition for identification requires that the network must include at least two students separated by a path of distance three, which is slightly stronger than the condition in BramoulleDjebbariFortin2009.
We illustrate these points through both a Monte Carlo simulation study and an empirical application using data from the National Longitudinal Study of Adolescent to Adult Health (Add Health). Estimating peer effects through our proposed method suggests that increasing the average GPA of peers by one point leads to a 0.856 point increase in students’ GPA. In contrast, the standard linear-in-means model using GPA as a proxy for student academic effort estimates this effect at 0.507. This substantial difference highlights that failing to account for two types of shocks at the school level may yield substantially biased estimates. Importantly, we also find that other plausible approaches, such as estimating a model that accounts for a single school-level fixed effect and incorporating a dummy variable for isolated students, fail to rectify the bias, as does excluding isolated students from the sample. Finally, we show that the standard model is sensitive to plausible forms of measurement error: when a share of students is misclassified as isolated, its peer effect estimate declines, whereas our double fixed effects model remains stable.
Isolated nodes—students without direct friendship connections—are a common feature in studies examining directed social networks. In the Add Health dataset, 22% of students have no friends, although 11% are not entirely isolated, as they are nominated by others as friends. Several studies using social network data from educational settings, such as alan2021building, conti2013popularity, boucher2021ethnic, banerjee2024changes, report similar patterns. We show that even with a small share of isolated nodes, the bias can be large.
We also extend our analysis to the case of endogenous networks. Network endogeneity can occur because we do not observe certain student characteristics, such as intellectual quotient (IQ), that may influence both students’ likelihood to form links with others and their GPA. We control for these unobserved characteristics using a two-stage estimation approach. Our method is nonparametric as in johnsson2021estimation. We do not impose a specific parametric restriction in the relationship between the unobserved characteristics and GPA. Our approach is similar to generalized additive models (GAM), which are widely employed in the nonparametric regression literature (see hastie2017generalized).
Our structural model and econometric approach can also be applied to study peer influence on other outcomes that depend on exerted effort. An example is the body mass index (BMI), which cannot be directly chosen (e.g., fy2015). People need to exert effort, such as developing healthy diet habits or engaging in physical exercise to improve their BMI. Peer influences are more related to effort than BMI. Another example is peer effects on a worker’s effort (e.g., mas2009peers; cornelissen2017peer). The observed outcome is generally worker’s productivity, whereas peer effects take source in effort.
Related Literature
There is a large literature studying social interactions both theoretically and empirically (durlauf2010social; blume2011identification). We follow the games in networks approach to the analysis of social interactions (see jackson2015games, for a comprehensive overview of this literature). BallesterZenou2006 analyze a noncooperative game with linear-quadratic utilities and strategic complementarities, in which each player decides how much effort they exert. Applying a similar framework to an education application, cpz2009 use GPA as a proxy for the exerted effort. We contribute to this strand of the literature by explicitly modelling the production function that captures how effort, along with other factors, translates into GPA, while distinguishing between unobserved shocks that directly impact GPA without affecting effort, from shocks that impact both academic effort and GPA. This leads to a reduced-form equation for GPA that differs from the standard linear-in-means peer effects specification.
This paper makes a methodological contribution to the extensive empirical literature on peer effects on educational outcomes (sacerdote2011peer; epple2011peer). We show both analytically and through an empirical application using AddHealth data that approximating student effort with GPA may result in biased estimates of peer effects, when some students in the network lack friends. Since isolated students are a common feature in many social network datasets, this finding highlights the limitations of using proxy variables, such as GPA, for estimating peer effects reliably.
Our paper also contributes to the econometric literature on peer effects (de2017econometrics; kline2020econometric). A key challenge in this field is the reflection problem (m1993). A recent wave of papers have addressed this issue by imposing conditions on the network structure (BramoulleDjebbariFortin2009; de2010identification). [^3] Other studies address the reflection problem using group size variation (davezies2009identification; lee2007identification) or imposing restrictions on the error terms (graham2008identifying; rose2017identification). For an overview of this literature, see bramoulle2020peer. Our contribution is to study the reflection problem in a setting where the GPA is influenced by various types of common shocks at the school level, and where students may have no peers. As in BramoulleDjebbariFortin2009, our main identification condition involves the network structure and can be readily tested in empirical applications.
The remainder of the paper is organized as follows. Section 2 presents the microeconomic foundations of the model using a network game in which students decide their academic effort. Section 3 describes the econometric model and addresses the identification and estimation of the parameters. Section 4 presents our empirical analysis using Add Health data. Section LABEL:sec:endo provides an extension of our framework to endogenous networks. Section LABEL:sec:conc concludes this paper.
2  The Structural Model
In this section, we introduce a structural model based on a game of complete information, where students choose their effort level, which impacts their educational achievements (GPA). Students’ effort levels are also influenced by the effort exerted by their peers. This model can also be applied to investigating peer effects on other inputs that are not typically directly observed by the econometrician. For instance, it can be used to examine peer effects on (unobserved) effort in a workplace setting where observed productivity serves as the outcome. Another application is peer effects on effort directed toward improving BMI, such as adopting healthy diet habits or engaging in physical exercise, which is more readily observable than the effort itself.
We consider $S$ independent schools and denote by $n_{s}$ the number of students in the $s$-th school, $s\in\{1,\dots,S\}$. Let $n$ be the total number of students, that is, $n=\sum_{s=1}^{S}n_{s}$. Each student $i$ in school $s$ has a GPA denoted by $y_{s,i}$ and observable characteristics represented by a $K$-vector $\mathbf{x}_{s,i}$. Students interact in their school through a directed network that can be represented by an $n_{s}\times n_{s}$ adjacency matrix $\mathbf{A}_{s}=[a_{s,ij}]_{ij}$, where $a_{s,ij}$ = 1 if student $j$ is $i^{\prime}s$ friend and $a_{s,ij}=0$ otherwise. We assume that $a_{s,ii}$ = 0 for all $i$ and $s$ so that students cannot interact with themselves. In addition, we only consider within-school interactions; students do not interact with peers in other schools (see cpz2009). We define the social interaction matrix $\mathbf{G}_{s}=[g_{s,ij}]_{ij}$ as the row-normalized adjacent matrix $\mathbf{A}_{s}$, that is, $g_{s,ij}=1/n_{s,i}$ if $j$ is a $i$’s friend and $g_{s,ij}=0$ otherwise, where $n_{s,i}$ is the number of friends of student $i$ within school $s$.
Student $i$ chooses their effort $e_{s,i}$, which in turn affects their GPA. More precisely, GPA is a function of student effort $e_{s,i}$ and a random term $\eta_{s,i}$ that captures unobservable characteristics. [^4] Whether or not the random term $\eta_{s,i}$ is observed by student $i$ or their peers is inconsequential for the analysis of the game. Following fruehwirth2013identifying and BoucherFortin2016, we posit that this relationship is given by the production function:
$y_{s,i}=\alpha_{s}+e_{s,i}+\eta_{s,i},$ (1)
where $\alpha_{s}$ captures unobserved school-level GPA shifters, such as teacher quality, operating as fixed effects. The linear production function is somewhat restrictive and can be generalized to a nonlinear or nonparametric production function (e.g., see fruehwirth2013identifying). However, identification of the resulting econometric model can be intractable, especially when the model allows for unobserved school heterogeneity. We also consider the production function $y_{s,i}=\alpha_{s}+\delta e_{s,i}+\mathbf{x}^{\prime}_{s,i}\boldsymbol{\theta}+\eta_{s,i}$, where $\delta>0$ and $\boldsymbol{\theta}$ is a vector of parameters. In Appendix LABEL:append:GPA, we show that both specifications imply the same reduced-form structure for the achievement equation we estimate below, although in the second case, the coefficients on own and contextual covariates should be interpreted as reduced-form composites. We discuss identification in Section 3.1. For parsimony, we retain (1) in the main text and present the alternative specification in Appendix LABEL:append:GPA.
The effort exerted and the GPA obtained provide students with a benefit that is captured by a payoff function, which, as in cpz2009 and BoucherFortin2016, takes a linear-quadratic form:
$u_{s,i}(e_{s,i},\mathbf{e}_{s,-i},y_{s,i})=\underbrace{(c_{s}+\mathbf{x}_{s,i}^{\prime}\boldsymbol{\beta}+\mathbf{g}_{s,i}\mathbf{X}_{s}\boldsymbol{\gamma}+\varepsilon_{s,i})y_{s,i}-\frac{e_{s,i}^{2}}{2}}_{\text{private sub-payoff}}+\underbrace{\lambda e_{s,i}\mathbf{g}_{s,i}\mathbf{e}_{s}}_{\text{social sub-payoff}},$ (2)
where $\mathbf{X}_{s}=(\mathbf{x}_{s,1},\dots,\mathbf{x}_{s,n_{s}})^{\prime}$, $\mathbf{g}_{s,i}$ is the $i$-th row of $\mathbf{G}_{s}$, $\mathbf{e}_{s,-i}=(e_{s,1},\dots,e_{s,i-1},e_{s,i+1},\dots,e_{s,n})^{\prime}$, $\mathbf{e}_{s}=(e_{s,1},\dots,e_{s,n_{s}})^{\prime}$, the term $\mathbf{g}_{s,i}\mathbf{e}_{s}$ is the average effort of peers, and $c_{s}$, $\boldsymbol{\beta}$, $\boldsymbol{\gamma}$ are unknown parameters. The parameter $\lambda$ captures endogenous peer effects. The payoff function (2) encompasses two components: a private sub-payoff and a social sub-payoff. The term $(c_{s}+\mathbf{x}_{s,i}^{\prime}\boldsymbol{\beta}+\mathbf{g}_{s,i}\mathbf{X}_{s}\boldsymbol{\gamma}+\varepsilon_{s,i})$ represents the benefit enjoyed per unit of GPA achieved, where $\varepsilon_{s,i}$ is the student type (observable by all students). This benefit accounts for student observed heterogeneity, as it depends on $\mathbf{x}_{s,i}$ and peer group average characteristics $\mathbf{g}_{s,i}\mathbf{X}_{s}$ termed contextual variables (see m1993). The benefit also accounts for school unobserved heterogeneity through the parameter $c_{s}$. The second term of the private sub-payoff, $e_{s,i}^{2}/2$ reflects the cost of exerting effort. The social sub-payoff $\lambda e_{s,i}\mathbf{g}_{s,i}\mathbf{e}_{s}$ implies that an increase in the average peer group’s effort $\mathbf{g}_{s,i}\mathbf{e}_{s}$ influences student $i$’s marginal payoff if $\lambda\neq 0$. When $\lambda>0$, the payoff function (2) implies complementarity between students’ and peers’ efforts, whereas $\lambda<0$ indicates substitutability in efforts. [^5] An alternative specification considers the social-payoff as $\frac{\lambda}{2}(e_{s,i}-\mathbf{g}_{s,i}\mathbf{e}_{s})^{2}$, which represents a social cost depending on the gap between the student’s effort level and the average peer effort. This specification leads to conformist preferences when $\lambda>0$. Our approach can also be extended to this alternative specification.
The parameters $\alpha_{s}$ and $c_{s}$ capture different unobserved shocks at the school level, and are conceptually different in terms of their policy implications. $\alpha_{s}$ captures unobserved shocks on GPA that do not affect student effort. These shocks, such as variation in teaching quality and school management, directly impact GPA irrespective of student effort. [^6] This reasoning holds because GPA is unbounded. An increase in $\alpha_{s}$ necessarily results in a higher GPA, and thus a higher payoff, regardless of the effort level. If we were to consider a framework where GPA is bounded, an increase in $\alpha_{s}$ may decrease the effort for students nearing the upper bound of GPA (e.g., see fruehwirth2013identifying). We address this concern directly by re-estimating our main specifications using a Tobit model in Appendix LABEL:append:app (Table LABEL:tab:estexotobit). The ordering of peer-effects estimates across models is unchanged. On the other hand, $c_{s}$ captures shocks on student preferences, particularly on the marginal payoff. For instance, interventions aimed at making students more aware of the returns to academic achievement could influence the marginal payoff. Such a shock can be captured by $c_{s}$ and would influence effort and consequently GPA through equation (1). We show that $\alpha_{s}$ and $c_{s}$ do not impact GPA in the same way (see Section 3.1).
By substituting $y_{s,i}$ from Equation (1) into Equation (2), we obtain a payoff function that does not depend on GPA (see Appendix LABEL:append:NE). This new payoff function defines a static game with complete information, in which students simultaneously choose their effort levels to maximize their payoff. The best response function for the students is given by:
$e_{s,i}=c_{s}+\lambda\mathbf{g}_{s,i}\mathbf{e}_{s}+\mathbf{x}_{s,i}^{\prime}\boldsymbol{\beta}+\mathbf{g}_{s,i}\mathbf{X}_{s}\boldsymbol{\gamma}+\varepsilon_{s,i}.$ (3)
In Equation (3), students’ levels of effort are expressed as a function of the average effort of their peers $\mathbf{g}_{s,i}\mathbf{e}_{s}$, observed students’ characteristics $\mathbf{x}_{s,i}$, and the average characteristics of peers $\mathbf{g}_{s,i}\mathbf{X}_{s}$ (contextual variables). The parameter $\lambda$ represents the impact of peers on a student’s effort level. A positive value of $\lambda$ indicates that a student’s effort level increases if their peers put in more effort. Furthermore, Equation (3) shows that the optimal effort level (and thus the resulting GPA) is influenced by shocks at the school level (on $c_{s}$) that affect student preferences. However, shocks directly affecting GPA through the parameter $\alpha_{s}$ do not impact effort.
The best response function (3) in matrix form can be expressed as $\mathbf{e}_{s}=c_{s}\mathbf{1}_{n_{s}}+\lambda\mathbf{G}_{s}\mathbf{e}_{s}+\mathbf{X}_{s}\boldsymbol{\beta}+\mathbf{G}_{s}\mathbf{X}_{s}\boldsymbol{\gamma}+\boldsymbol{\varepsilon}_{s}$, where $\boldsymbol{\varepsilon}_{s}=(\varepsilon_{s,1},\dots,\varepsilon_{s,n_{s}})^{\prime}$ and $\mathbf{1}_{n_{s}}$ is an $n_{s}$-vector of ones. A solution of this equation in $\mathbf{e}_{s}$ is a Nash equilibrium (NE) of the game. As $\mathbf{G}_{s}$ is row-normalized, the NE is unique under Assumption 2.1 and can be expressed as $\mathbf{e}_{s}=(\mathbf{I}_{n_{s}}-\lambda\mathbf{G}_{s})^{-1}(c_{s}\mathbf{1}_{n_{s}}+\mathbf{X}_{s}\boldsymbol{\beta}+\mathbf{G}_{s}\mathbf{X}_{s}\boldsymbol{\gamma}+\boldsymbol{\varepsilon}_{s})$, where $\mathbf{I}_{n_{s}}$ is the $n_{s}\times n_{s}$ identity matrix (see Appendix LABEL:append:NE).
Assumption 2.1.
$|\lambda|<1$.
The condition $|\lambda|<1$ implies that students do not increase (in absolute value) their effort as much as the increase in the effort of their peers. Put differently, when the average effort of a student’s friends increases by one unit, the corresponding change in the student’s own effort is less than one unit in absolute value.
3  The Econometric Model and Identification Strategy
If effort were directly observable, we could estimate the peer effect parameter from Equation (3) following kelejian1998generalized. However, as we do not observe effort, the equation cannot be estimated directly. Instead, we derive from this equation an estimable equation based on GPA, which is influenced by effort. We show that this equation is econometrically different from the equation that we obtain if we proxy effort using GPA in Equation (2). We also present an identification strategy to identify the peer effect parameter $\lambda$ in the effort equation.
3.1 Reduced-Form Equation for GPA
From Equation (1), we express effort as a function of GPA and replace this expression in Equation (3). This yields a reduced-form equation for GPA that does not directly depend on effort (see Appendix LABEL:append:GPA). The equation is given by
$y_{s,i}=\kappa_{s,i}+\lambda\mathbf{g}_{s,i}\mathbf{y}_{s}+\mathbf{x}^{\prime}_{s,i}\boldsymbol{\beta}+\mathbf{g}_{s,i}\mathbf{X}_{s}\boldsymbol{\gamma}+\eta_{s,i}-\lambda\mathbf{g}_{s,i}\boldsymbol{\eta}_{s}+\varepsilon_{s,i},$ (4)
where $\kappa_{s,i}=c_{s}+(1-\lambda\mathbf{g}_{s,i}\mathbf{1}_{n_{s}})\alpha_{s}$.
If instead, we proxy effort by GPA in the payoff function (2), the resulting reduced-form equation of GPA would be:
$y_{s,i}=\bar{c}_{s}+\bar{\lambda}\mathbf{g}_{s,i}\mathbf{y}_{s}+\mathbf{x}_{s,i}^{\prime}\boldsymbol{\bar{\beta}}+\mathbf{g}_{s,i}\mathbf{X}_{s}\boldsymbol{\bar{\gamma}}+\bar{\varepsilon}_{s,i},$ (5)
where $\bar{\lambda}$ is the peer effect parameter, $\boldsymbol{\bar{\beta}}$ and $\boldsymbol{\bar{\gamma}}$ measure the effects of own and contextual characteristics, $\bar{c}_{s}$ controls for unobserved school heterogeneity and $\bar{\varepsilon}_{s,i}$ is an error term. We will refer to this specification as the standard (classical) model.
Let $\mathcal{V}^{NI}_{s}$ denote the subsample of students of school $s$ who have friends (non-isolated) and $\mathcal{V}^{I}_{s}$ denote the subsample of students of school $s$ who have no friends (isolated). As $\mathbf{G}_{s}$ is row-normalized, we have $\mathbf{g}_{s,i}\mathbf{1}_{n_{s}}=1$ if $i\in\mathcal{V}^{NI}_{s}$ and $\mathbf{g}_{s,i}\mathbf{1}_{n_{s}}=0$ otherwise. Thus, $\kappa_{s,i}=\kappa^{I}_{s}$ if $i\in\mathcal{V}^{I}_{s}$ and $\kappa_{s,i}=\kappa^{NI}_{s}$ if $i\in\mathcal{V}^{NI}_{s}$, where $\kappa^{I}_{s}=c_{s}+\alpha_{s}$ and $\kappa^{NI}_{s}=c_{s}+(1-\lambda)\alpha_{s}$. If there are no isolated students in the network, $\kappa_{s,i}$ would be a simple school fixed effect, and our framework would be equivalent to the classical model. [^7] An isolated student is a student who has no friends. However, this student may be a friend of others. We later refer to a fully isolated student as a student who has no friends and who is not a friend of others. The difference between the standard model and our framework is that the standard model has a single intercept term per school. In Equation (4), $\kappa_{s,i}$ accounts for unobserved school-level heterogeneity depending on whether the student $i$ has friends or not. We would have $\kappa^{NI}_{s}=\kappa^{I}_{s}$ if and only if $\alpha_{s}=0$, $\lambda=0$, or if every student has friends. These conditions may not hold in many settings.
To appropriately account for the unobserved factor $\kappa_{s,i}$ in Equation (4), the above discussion suggests that we need to incorporate both school-fixed effects and a school-specific variable indicating whether a student has friends or not. This involves including $S$ school dummy variables and $S$ dummy variables indicating whether each student has friends or not. Each of the latter dummy variables corresponds to one school and takes the value of one if the student has friends. The rationale for this lies in the distinction between $\alpha_{s}$ and $c_{s}$, which affect GPA differently. To understand why distinguishing between these two shocks requires controlling for whether a student is isolated or not, it is important to consider the implications of the two shocks. As per Equations (3) and (4), an increase in $\alpha_{s}$ (e.g., by improving teaching quality) suggests an increase in GPA without affecting effort. Importantly, this increase does not depend on whether the student is isolated and does not involve a social multiplier effect. [^8] Equation (4) implies that the variation in $\mathbf{y}_{s}$, denoted $\Delta^{\alpha}\mathbf{y}_{s}$, following an increase $\Delta\alpha_{s}$ in $\alpha_{s}$ is such that $\Delta^{\alpha}\mathbf{y}_{s}=\Delta\alpha_{s}(\mathbf{I}_{n_{s}}-\lambda\mathbf{G}_{s})\mathbf{1}_{n_{s}}+\lambda\mathbf{G}_{s}\Delta^{\alpha}\mathbf{y}_{s}$. This implies that $(\mathbf{I}_{n_{s}}-\lambda\mathbf{G}_{s})\Delta^{\alpha}\mathbf{y}_{s}=\Delta\alpha_{s}(\mathbf{I}_{n_{s}}-\lambda\mathbf{G}_{s})\mathbf{1}_{n_{s}}$ and thus, $\Delta^{\alpha}\mathbf{y}_{s}=\Delta\alpha_{s}\mathbf{1}_{n_{s}}$. Hence, the increase in $\alpha_{s}$ results in the same increase in GPA for all students. In contrast, a preference shock on $c_{s}$ could result in a social multiplier effect on the effort (according to Equation (3)), and thereby on GPA. This social multiplier effect is only present among students who have friends. Therefore, the distinction between the two types of shocks is essentially captured by the student’s social status: unlike for $\alpha_{s}$, the impact of $c_{s}$ on GPA is contingent upon whether the student is isolated or not.
The main distinction between the standard model and Equation (4) arises from an omitted-variable problem in the standard model. This omitted variable is $\lambda\alpha_{s}\mathbf{g}_{s,i}\mathbf{1}_{n_{s}}$ and may introduce a discrepancy between the peer-effect estimates of the two models. Since this omitted variable depends on the network matrix $\mathbf{G}_{s}$, it can induce correlation between the peer regressor $\mathbf{G}_{s}\mathbf{y}_{s}$ and the structural error term. Consequently, network-based instruments constructed using $(\mathbf{G}_{s},\mathbf{X}_{s})$ (see BramoulleDjebbariFortin2009) may violate the exclusion restriction. The sign of the bias in the peer effect parameter depends on the sign of the covariance between the instrument and the omitted variable $\lambda\alpha_{s}\mathbf{g}_{s,i}\mathbf{1}_{n_{s}}$, as well as on the sign of the covariance between the instrument and the instrumented variable $\mathbf{G}_{s}\mathbf{y}_{s}$. Determining this sign is not straightforward because $\mathbf{X}_{s}$ includes many variables, each potentially leading to an instrument that influences the omitted variable and $\mathbf{G}_{s}\mathbf{y}_{s}$ differently.
3.2 Identification and Estimation
Let $n^{NI}_{s}$ and $n^{I}_{s}$ denote the number of students in school $s$ with and without peers, respectively. In addition, let $\boldsymbol{\ell}^{NI}_{s}=\mathbf{G}_{s}\mathbf{1}_{n_{s}}$ and $\boldsymbol{\ell}^{I}_{s}=\mathbf{1}_{n_{s}}-\boldsymbol{\ell}^{NI}_{s}$. Equation (4) can be written in matrix form at the school level as
$\mathbf{y}_{s}=\kappa^{I}_{s}\boldsymbol{\ell}^{I}_{s}+\kappa^{NI}_{s}\boldsymbol{\ell}^{NI}_{s}+\lambda\mathbf{G}_{s}\mathbf{y}_{s}+\mathbf{X}_{s}\boldsymbol{\beta}+\mathbf{G}_{s}\mathbf{X}_{s}\boldsymbol{\gamma}+(\mathbf{I}_{n_{s}}-\lambda\mathbf{G}_{s})\boldsymbol{\eta}_{s}+\boldsymbol{\varepsilon}_{s}.$ (6)
Note that the number of unknown parameters to be estimated in Equation (6) grows to infinity with the number of schools (there are $2S$ dummy variables). This issue is known as an incidental parameter problem and may lead to inconsistent estimators. A common approach to consistently estimate the model is to eliminate the fixed effects $\kappa^{I}_{s}$ and $\kappa^{NI}_{s}$. To do so, we define $\mathbf{J}_{s}:=\mathbf{I}_{n_{s}}-\dfrac{1}{n^{I}_{s}}\boldsymbol{\ell}^{I}_{s}\boldsymbol{\ell}^{I\prime}_{s}-\dfrac{1}{n^{NI}_{s}}\boldsymbol{\ell}^{NI}_{s}\boldsymbol{\ell}^{NI\prime}_{s}$ and impose by convention that $\dfrac{1}{n^{I}_{s}}\boldsymbol{\ell}^{I}_{s}\boldsymbol{\ell}^{I\prime}_{s}=0$ if $n^{I}_{s}=0$, and that $\dfrac{1}{n^{NI}_{s}}\boldsymbol{\ell}^{NI}_{s}\boldsymbol{\ell}^{NI\prime}_{s}=0$ if $n^{NI}_{s}=0$. Note that $\boldsymbol{\ell}^{NI\prime}_{s}\boldsymbol{\ell}^{NI}_{s}=n^{NI}_{s}$, $\boldsymbol{\ell}^{I\prime}_{s}\boldsymbol{\ell}^{I}_{s}=n^{I}_{s}$, $\boldsymbol{\ell}^{NI\prime}_{s}\boldsymbol{\ell}^{I}_{s}=0$, $\boldsymbol{\ell}^{I\prime}_{s}\boldsymbol{\ell}^{NI}_{s}=0$. Thus, $\mathbf{J}_{s}\boldsymbol{\ell}^{I}_{s}=\mathbf{J}_{s}\boldsymbol{\ell}^{NI}_{s}=0$. One can eliminate the term $\kappa^{I}_{s}\boldsymbol{\ell}^{I}_{s}+\kappa^{NI}_{s}\boldsymbol{\ell}^{NI}_{s}$ by premultiplying each term of Equation (6) by the matrix $\mathbf{J}_{s}$. [^9] By premultiplying each term by $\mathbf{J}_{s}$, we consider Equation (6) in deviation to the average within the student group, that is, $\mathcal{V}^{I}_{s}$ or $\mathcal{V}^{I}_{s}$. This eliminates the parameters $\kappa^{I}_{s}$ and $\kappa^{NI}_{s}$. This implies that
$\mathbf{J}_{s}\mathbf{y}_{s}=\lambda\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{y}_{s}+\mathbf{J}_{s}\mathbf{X}_{s}\boldsymbol{\beta}+\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{X}_{s}\boldsymbol{\gamma}+\mathbf{J}_{s}(\mathbf{I}_{n_{s}}-\lambda\mathbf{G}_{s})\boldsymbol{\eta}_{s}+\mathbf{J}_{s}\boldsymbol{\varepsilon}_{s}.$ (7)
The random term $\boldsymbol{v}_{s}:=\mathbf{J}_{s}(\mathbf{I}_{n_{s}}-\lambda\mathbf{G}_{s})\boldsymbol{\eta}_{s}+\mathbf{J}_{s}\boldsymbol{\varepsilon}_{s}$ is comprised of two error vectors $\boldsymbol{\eta}_{s}$ and $\boldsymbol{\varepsilon}_{s}$. To consistently estimate $\boldsymbol{\psi}:=(\lambda,\penalty 10000\ \boldsymbol{\beta}^{\prime},\penalty 10000\ \boldsymbol{\gamma}^{\prime})^{\prime}$, we impose the following assumption.
Assumption 3.1.
For any $s=1,\dots,S$, $\mathbb{E}(\boldsymbol{\eta}_{s}|\mathbf{G}_{s},\mathbf{X}_{s})=0$ and $\mathbb{E}(\boldsymbol{\varepsilon}_{s}|\mathbf{G}_{s},\mathbf{X}_{s})=0$
Assumption 3.1 implies that $\mathbf{X}_{s}$ and $\mathbf{G}_{s}$ are exogenous with respect to $\boldsymbol{\eta}_{s}$ and $\boldsymbol{\varepsilon}_{s}$. This suggests that there is no omission of important variables in $\mathbf{X}_{s}$, which are captured by $\boldsymbol{\eta}_{s}$ and $\boldsymbol{\varepsilon}_{s}$. We later relax this assumption in our approach to controlling for network endogeneity by allowing for $\mathbf{X}_{s}$ and $\mathbf{G}_{s}$ to depend on $\boldsymbol{\eta}_{s}$ and $\boldsymbol{\varepsilon}_{s}$ through unobserved factors to the econometrician (see Section LABEL:sec:endo).
Identification and Estimation of $\boldsymbol{\psi}$
Identification in peer effects models can be challenging, particularly due to the reflection problem (m1993). BramoulleDjebbariFortin2009 address this problem and provide necessary and sufficient conditions for identification. Their main condition requires that $\mathbf{I}_{n_{s}}$, $\mathbf{G}_{s}$, $\mathbf{G}_{s}^{2}$, and $\mathbf{G}_{s}^{3}$ are linearly independent. However, their approach does not apply to our framework because their identification results assume that there are no isolated students (see their Section 3.1).101010bramoulle2020peer also discuss the case involving isolated students. They argue that the presence of isolated students can help identify the peer effect parameter (see their Section 2.1.1). Recall that the two intercepts of our model are $\kappa^{I}_{s}=c_{s}+\alpha_{s}$ and $\kappa^{NI}_{s}=c_{s}+(1-\lambda)\alpha_{s}$. The fact that only one intercept depends on $\lambda$ implies that $\lambda$ can be identified using isolated students if $\kappa^{I}_{s}=\kappa^{NI}_{s}$. However, as we allow GPA shocks $\alpha_{s}$ to be different from preference shocks $c_{s}$ in terms of their impact on effort, this condition is unlikely to hold. Consequently, the presence of isolated students does not help identify $\lambda$ in our case.
Given that existing identification results do not directly apply to our model, we extend the analysis in BramoulleDjebbariFortin2009. We derive easy-to-verify conditions to address the reflection problem when the network includes students without friends.
Assumption 3.2.
(i) $\lambda\boldsymbol{\beta}+\boldsymbol{\gamma}\neq 0$; (ii) There are students in the network separated by a link of distance three.
Condition (i) is equivalent to stating that GPA is influenced by at least one contextual variable. [^11] See Equation LABEL:eq:append:Dy, which quantifies the total effect of an increase in a contextual variable on GPA. With several characteristics in $\mathbf{X}_{s}$, this condition can be satisfied. Condition (ii) is slightly stronger than the assumption that $\mathbf{I}_{n_{s}}$, $\mathbf{G}_{s}$, $\mathbf{G}_{s}^{2}$, and $\mathbf{G}_{s}^{3}$ are linearly independent. It means that there are students who have connections to students that extend to three degrees of separation—friends of friends of friends—who are neither directly their friends nor friends of their friends (see an illustration in Figure 1 below).
Under Assumption 3.2, we show that $\boldsymbol{\psi}$ is identified. Before formally presenting this result and its proof, we first provide an illustration. We consider a network with a link of distance three as depicted in Figure 1: $i_{2}$ is at three nodes from $i_{1}$; they are a friend of a friend of a friend of $i_{1}$.
$i_{2}$$i_{4}$$i_{1}$$i_{3}$
Figure 1: Solving the reflection problem
Note: $\rightarrow$ means that the node on the right side is a friend of the node on the left side.
We employ a proof by contradiction. The nonidentification issue arises when $\mathbb{E}(\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{y}_{s}|\mathbf{G}_{s},\mathbf{X}_{s})$ is perfectly collinear with $\mathbf{J}_{s}\mathbf{X}_{s}$ and $\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{X}_{s}$. For a non-isolated student, this suggests that there exist vector of parameters, $\dot{\boldsymbol{\beta}}$ and $\dot{\boldsymbol{\gamma}}$, such that $\mathbb{E}(\mathbf{g}_{s,i}\mathbf{y}_{s}-y^{NI}_{s}|\mathbf{G}_{s},\mathbf{X}_{s})=(\mathbf{x}_{s,i}-\mathbf{\hat{x}}_{s})^{\prime}\dot{\boldsymbol{\beta}}+(\mathbf{g}_{s,i}\mathbf{X}_{s}-\bar{\mathbf{x}}^{NI}_{s})^{\prime}\dot{\boldsymbol{\gamma}}$, where $y^{NI}_{s}$, $\mathbf{x}^{NI}_{s}$, and $\bar{\mathbf{x}}^{NI}_{s}$ are respectively the averages of $\mathbf{g}_{s,i}\mathbf{y}_{s}$, $\mathbf{x}_{s,i}$, and $\mathbf{g}_{s,i}\mathbf{X}_{s}$ in $\mathcal{V}^{NI}_{s}$. The variables $\mathbf{g}_{s,i}\mathbf{y}_{s}$, $\mathbf{x}_{s,i}$, and $\mathbf{g}_{s,i}\mathbf{X}_{s}$ are taken in deviation with respect to their average in $\mathcal{V}^{NI}_{s}$ because of the matrix $\mathbf{J}_{s}$ that multiplies the terms of Equation (7). If we take the previous equation in difference between students $i_{1}$ and $i_{3}$, we obtain
$\mathbb{E}(\mathbf{g}_{s,i_{1}}\mathbf{y}_{s}-\mathbf{g}_{s,i_{3}}\mathbf{y}_{s}|\mathbf{G}_{s},\mathbf{X}_{s})=(\mathbf{x}_{s,i_{1}}-\mathbf{x}_{s,i_{3}})^{\prime}\dot{\boldsymbol{\beta}}+(\mathbf{g}_{s,i_{1}}\mathbf{X}_{s}-\mathbf{g}_{s,i_{3}}\mathbf{X}_{s})^{\prime}\dot{\boldsymbol{\gamma}}.$ (8)
We now show that Condition (i) of Assumption 3.2 is not compatible with Equation (8). As $i_{3}$ is $i_{1}$’s only friend and $i_{4}$ is $i_{3}$’s only friend, we have $\mathbf{g}_{s,i_{1}}\mathbf{y}_{s}=y_{s,i_{3}}$, $\mathbf{g}_{s,i_{3}}\mathbf{y}_{s}=y_{s,i_{4}}$, $\mathbf{g}_{s,i_{1}}\mathbf{X}_{s}=\mathbf{x}_{s,i_{3}}$, and $\mathbf{g}_{s,i_{3}}\mathbf{X}_{s}=\mathbf{x}_{s,i_{4}}$. Thus, Equation (8) implies $y_{i_{3}}^{e}-y_{i_{4}}^{e}=(\mathbf{x}_{s,i_{1}}-\mathbf{x}_{s,i_{3}})^{\prime}\dot{\boldsymbol{\beta}}+(\mathbf{x}_{s,i_{3}}-\mathbf{x}_{s,i_{4}})^{\prime}\dot{\boldsymbol{\gamma}}$, where $y_{i}^{e}=\mathbb{E}(y_{s,i}|\mathbf{G}_{s},\mathbf{X}_{s})$. Since $\mathbf{x}_{s,i_{2}}$ does not appear in the previous equation, it follows that, an increase in $\mathbf{x}_{s,i_{2}}$ (keeping $\mathbf{x}_{s,i_{1}}$, $\mathbf{x}_{s,i_{3}}$, and $\mathbf{x}_{s,i_{4}}$ fixed) has no impact on $y_{i_{3}}^{e}-y_{i_{4}}^{e}$. Put differently, this increase has either no impact on $y_{i_{3}}^{e}$ and $y_{i_{4}}^{e}$ or the same impact on both $y_{i_{3}}^{e}$ and $y_{i_{4}}^{e}$. The first implication is not possible because $\mathbf{x}_{s,i_{2}}$ is the vector of contextual variables for $i_{4}$ and Condition (i) of Assumption 3.2 implies that GPA is influenced by at least one contextual variable. Moreover, $\mathbf{x}_{s,i_{2}}$ cannot influence $y_{i_{3}}^{e}$ and $y_{i_{4}}^{e}$ in the same way because $i_{2}$ is a direct friend of $i_{4}$ and not directly linked to $i_{3}$. Consequently, $\mathbb{E}(\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{y}_{s}|\mathbf{G}_{s},\mathbf{X}_{s})$ cannot be a linear combination of $\mathbf{J}_{s}\mathbf{X}_{s}$ and $\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{X}_{s}$.
As in the case of the standard model, the regressor $\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{y}_{s}$ is endogenous in Equation (7). However, it can be instrumented by $\mathbf{J}_{s}\mathbf{G}_{s}^{2}\mathbf{X}_{s}$ because $\mathbf{J}_{s}\mathbf{G}_{s}^{2}\mathbf{X}_{s}$ is an excluded regressor with is correlated to $\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{y}_{s}$(BramoulleDjebbariFortin2009). Let $\boldsymbol{\hat{\psi}}$ be the IV estimator of $\boldsymbol{\psi}$ using the instrument vector $\mathbf{J}_{s}\mathbf{G}_{s}^{2}\mathbf{X}_{s}$. [^12] To avoid a weak instrument issue, the pool of instruments can also be expanded to $\mathbf{J}_{s}[\mathbf{G}_{s}^{2}\mathbf{X}_{s},\penalty 10000\ \dots,\penalty 10000\ \mathbf{G}_{s}^{p}\mathbf{X}_{s}]$ for some integer $p>2$. We have the following result.
Proposition 3.1.
Under Assumptions 2.13.2 and LABEL:ass:app:ident:gmm (stated in Appendix LABEL:append:ident:reflection), $\boldsymbol{\psi}$ is globally identified, $\boldsymbol{\hat{\psi}}$ is a consistent estimator, and $)\displaystyle\sqrt{n}(\boldsymbol{\hat{\psi}}-\boldsymbol{\psi}_{0})\overset{d}{\to}\mathcal{N}\big(0,\lim_{n\to\infty}n\operatorname{\mathbb{V}}(\boldsymbol{\hat{\psi}}$$\big)$, where $\boldsymbol{\psi}_{0}$ is the true value of $\boldsymbol{\psi}$.
The proof of Proposition 3.1 is presented in Appendix LABEL:append:ident:reflection. The consistency and asymptotic normality of $\boldsymbol{\hat{\psi}}$ are directly derived from kelejian1998generalized.
Asymptotic variance of $\boldsymbol{\hat{\psi}}$
We now discuss how the asymptotic variance of $\boldsymbol{\hat{\psi}}$ can be consistently estimated. One simple approach is to use the robust estimator à la white1980heteroskedasticity that allows for the variance of the component of $\boldsymbol{\eta}_{s}$ and $\boldsymbol{\varepsilon}_{s}$ to vary across schools. Let $\mathbf{\tilde{X}}_{s}=\mathbf{J}_{s}[\mathbf{X}_{s},\penalty 10000\ \mathbf{G}_{s}\mathbf{X}_{s}]$, $\mathbf{R}_{s}=[\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{y}_{s},\penalty 10000\ \mathbf{\tilde{X}}_{s}]$, $\mathbf{Z}_{s}=[\mathbf{J}_{s}\mathbf{G}_{s}^{2}\mathbf{X}_{s},\penalty 10000\ \mathbf{\tilde{X}}_{s}]$, $\mathbf{R}^{\prime}\mathbf{Z}=\sum_{s=1}^{S}\mathbf{R}^{\prime}_{s}\mathbf{Z}_{s}$, $\mathbf{Z}^{\prime}\mathbf{Z}=\sum_{s=1}^{S}\mathbf{Z}^{\prime}_{s}\mathbf{Z}_{s}$. Let also $\hat{\boldsymbol{v}}_{s}:=\mathbf{J}_{s}\mathbf{y}_{s}-\mathbf{R}_{s}\boldsymbol{\hat{\psi}}$, that is, $\hat{\boldsymbol{v}}_{s}$ is the residual vector from Equation (7). We denote by $\operatorname{diag}$ the bloc diagonal matrix operator. The asymptotic variance of $\boldsymbol{\hat{\psi}}$ can be estimated by $\mathbf{\hat{B}}_{n}^{-1}\mathbf{\hat{D}}_{n}\mathbf{\hat{B}}_{n}^{-1}/n$, where $\mathbf{\hat{B}}_{n}=(\mathbf{R}^{\prime}\mathbf{Z})(\mathbf{Z}^{\prime}\mathbf{Z})^{-1}(\mathbf{R}^{\prime}\mathbf{Z})^{\prime}/n$ and $\mathbf{\hat{D}}_{n}=(\mathbf{R}^{\prime}\mathbf{Z})(\mathbf{Z}^{\prime}\mathbf{Z})^{-1}\operatorname{diag}\{\mathbf{Z}_{1}^{\prime}\hat{\boldsymbol{v}}_{1}\hat{\boldsymbol{v}}_{1}^{\prime}\mathbf{Z}_{1},\dots,\mathbf{Z}_{S}^{\prime}\hat{\boldsymbol{v}}_{S}\hat{\boldsymbol{v}}_{S}^{\prime}\mathbf{Z}_{S}\}(\mathbf{Z}^{\prime}\mathbf{Z})^{-1}(\mathbf{R}^{\prime}\mathbf{Z})^{\prime}/n$.
However, for this estimator of the asymptotic variance, we need to impose that $n_{s}$ is bounded, that is, the network is composed of many bounded and independent schools. This assumption is important for $\mathbf{Z}_{s}^{\prime}\hat{\boldsymbol{v}}_{s}$ to be $O_{p}(1)$ and have a second-order moment. We also present another estimator that does not require this condition. This estimator is based on the covariance structure of the error term $\boldsymbol{v}_{s}$. For this second approach, we set the following assumptions.
Assumption 3.3.
(i) $\mathbf{G}_{s}$is uniformly bounded in column sum; (ii) $\lambda\neq 0$; (iii) $(\eta_{s,i},\penalty 10000\ \varepsilon_{s,i})$is independently distributed across $i$ such that $\mathbb{E}(\eta_{s,i}^{2}|\mathbf{G}_{s},\mathbf{X}_{s})=\sigma_{\eta}^{2}>0$, $\mathbb{E}(\varepsilon_{s,i}^{2}|\mathbf{G}_{s},\mathbf{X}_{s})=\sigma_{\epsilon}^{2}>0$, $\mathbb{E}(\varepsilon_{s,i}\eta_{s,i}|\mathbf{G}_{s},\mathbf{X}_{s})=\rho\sigma_{\eta}\sigma_{\epsilon}$, and $\lvert\rho\rvert\leq 1$.
If both $S$ and $n_{s}$ tend to infinity, Condition (i) rules out the cases where the sum of certain columns of $\mathbf{G}_{s}$ is unbounded. This condition is also considered by (lee2004asymptotic) in the case of the standard model. Condition (iii) imposes constant variances for the error terms $\eta_{s,i}$ and $\varepsilon_{s,i}$ but accounts for their potential correlation. This is important as $\eta_{s,i}$ and $\varepsilon_{s,i}$ characterize the same student $i$. The parameter $\rho$ measures the correlation between $\eta_{s,i}$ and $\varepsilon_{s,i}$. We impose no restrictions on the joint distribution of $(\eta_{s,i},\penalty 10000\ \varepsilon_{s,i})$. The restriction $\lambda\neq 0$ of Condition (ii) is necessary for identifying $\sigma_{\eta}$, $\sigma_{\epsilon}$, and $\rho$. If $\lambda=0$, the disturbance of Equation (7) would be $\mathbf{J}_{s}\boldsymbol{\eta}_{s}+\mathbf{J}_{s}\boldsymbol{\varepsilon}_{s}$, and one cannot disentangle $\sigma_{\eta}$, $\sigma_{\epsilon}$, and $\rho$.
[…]
Proposition 3.2.
[…]
3.3 Simulation Study
[…]
We consider two exogenous variables in the matrix $\mathbf{X}_{s}$: $\mathbf{x}_{1s}\sim N(E_{1s},\penalty 10000\ 16)$ and $\mathbf{x}_{2s}\sim\text{Poisson}(E_{2s})$, where $E_{1s}$, $E_{2s}$ are fixed for each school and drawn from a Uniform distribution over [0, $0,\penalty 10000\ 1$0]. The distribution of these two control variables is specific to each school.
We investigate three data-generating processes (DGPs) denoted A, B, and C. For all DGPs, we assign the following values to the parameters: $\lambda=0.7$, $\boldsymbol{\beta}=(1,\penalty 10000\ 1.5)^{\prime}$, $\boldsymbol{\gamma}=(5,\penalty 10000\ -3)^{\prime}$, $\sigma^{2}_{\epsilon}=8$, $\sigma^{2}_{\eta}=15$, and $\rho=0.4$. We assume that $(\eta_{s,i},\penalty 10000\ \varepsilon_{s,i})$ follows a bivariate normal distribution. We consider the same school preference shock for all DGPs, defined as $c_{s}=-1.5q_{90}(\mathbf{x}_{2s})$, where $q_{a}(.)$ represents the $a$-th percentile. Generating $c_{s}$ in this way ensures that $c_{s}$ is a fixed effect because it is dependent on the control variables. To illustrate how the nature of the shock $\alpha_{s}$ can affect the results, we vary the definition of $\alpha_{s}$ for each DGP. For DGP A, we set $\alpha_{s}=0$ for all $s$. For DGP C, we define $\alpha_{s}=10q_{90}(\mathbf{x}_{1s})$. For DGP B, we set $\alpha_{s}$ to the average of the values of $\alpha_{1},\penalty 10000\ \dots,\penalty 10000\ \alpha_{S}$ obtained for DGP C. For both DGPs A and B, $\alpha_{s}$ does not vary across schools. The only difference is that $\alpha_{s}=0$ for DGP A, whereas $\alpha_{s}\neq 0$ for DGP B.
We consider three estimation approaches for each DGP. The most flexible estimation approach follows Equation (6), which can be expressed as follows:
$y_{s,i}=\kappa^{NI}_{s}\ell^{NI}_{s,i}+\kappa^{I}_{s}(1-\ell^{NI}_{s,i})+\lambda\mathbf{g}_{s,i}\mathbf{y}_{s}+\mathbf{x}^{\prime}_{s,i}\boldsymbol{\beta}+\mathbf{g}_{s,i}\mathbf{X}_{s}\boldsymbol{\gamma}+\tilde{v}_{s,i},$ (9)
where $\ell^{NI}_{s,i}$ is a dummy variable indicating whether the student is not isolated, $\kappa^{I}_{s}=c_{s}+\alpha_{s}$, $\kappa^{NI}_{s}=c_{s}+(1-\lambda)\alpha_{s}$, and $\tilde{v}_{s,i}=\eta_{s,i}-\lambda\mathbf{g}_{s,i}\boldsymbol{\eta}_{s}+\varepsilon_{s,i}$. This specification, referred to as Model 3, is our proposed model. Additionally, we estimate variants of this specification that impose constraints on $c_{s}$ and $\alpha_{s}$. These constrained specifications align with those commonly estimated in the literature when approximating students’ efforts by their GPA:
  • Model 1: Specification (9) with school fixed effects, which assumes that $\kappa^{I}_{s}=\kappa^{NI}_{s}=\kappa_{s}$, where $\kappa_{s}$ varies across schools.
  • Model 2: Specification (9) with one school fixed effect $\kappa_{s}$ shared by all students in school $s$ plus a single global indicator for being isolated with coefficient $\phi$. Equivalently, $\kappa_{s}^{NI}=\kappa_{s}$ and $\kappa_{s}^{I}=\kappa_{s}+\phi$.
Table 1 presents a summary of our Monte Carlo results across 1,000 replications.
Model 1 is the standard model commonly employed in the literature, especially when GPA approximates effort. This model demonstrates strong performance with DGP A, despite the presence of isolated students. Interestingly, this result suggests that the dummy variable for isolated students might not be relevant, even when the network includes students without friends. As this dummy variable accounts for GPA shocks without influencing effort, its absence in DGP A (because $\alpha_{s}=0$) does not compromise the validity of estimates. However, Model 1 leads to biased estimates for DGPs B and C, where $\alpha_{s}\neq 0$. Importantly, these biases carry substantial policy implications. Even with only 5% of isolated nodes, peer effect estimates are biased downward by over 35%. The bias seems not to be influenced by the share of isolated nodes.
Model 2 augments a single school fixed effect with a single global isolated-student dummy. Some researchers may adopt this specification to address the fact that isolated nodes have zero values for contextual variables. For instance, the share of female friends is zero for both isolated nodes and non-isolated nodes whose friends are exclusively male. To differentiate between these two cases, a dummy variable for isolated nodes can be incorporated into the model. While this approach corrects the bias observed in DGP B, it still results in a downward bias in DGP C. The bias is approximately 15% when the share of isolated nodes is 5%, increasing to 25% when 22% of nodes are isolated. This bias arises because of the assumption in DGP C that $\alpha_{s}$ varies across schools. [^13] As pointed out in Section 3.1, the bias depends on the sign of the covariance between the instrument and the omitted variable $\lambda\alpha_{s}\mathbf{g}_{s,i}\mathbf{1}_{n_{s}}$, as well as on the sign of the covariance between the instrument and the instrumented variable $\mathbf{G}_{s}\mathbf{y}_{s}$. It turns out that, by setting $\boldsymbol{\beta}=(-1,\penalty 10000\ -1.5)^{\prime}$ and $\boldsymbol{\gamma}=(-5,\penalty 10000\ 3)^{\prime}$—i.e., to the opposite of their actual values—so that the sign of the covariance between the instrument and $\mathbf{G}_{s}\mathbf{y}_{s}$ changes—Models 1 and 2 overestimate peer effects. We do not present these simulation results for brevity.
Model 3, which is our preferred specification (Equation (9)), performs well across all DGPs. Yet, it is worth noting that the standard deviations of the estimates for DGPs A and B are higher in Model 3 compared to Model 2. This is because both models are suitable for these DGPs, whereas Model 3 is overparameterized when the true isolation gap is zero or constant.
Table 1: Simulation results
Model 1 Model 2 Model 3
Mean Sd Mean Sd Mean Sd
5% of Isolated Nodes
DGP A
$\lambda=0.7$ $0.700$ $0.015$ $0.700$ $0.015$ $0.701$ $0.022$
$\beta_{0,1}=1$ $1.000$ $0.047$ $1.000$ $0.047$ $1.000$ $0.048$
$\beta_{0,2}=1.5$ $1.500$ $0.084$ $1.000$ $0.047$ $1.500$ $0.084$
$\gamma_{0,1}=5$ $4.996$ $0.091$ $4.996$ $0.092$ $4.997$ $0.092$
$\gamma_{0,2}=-3$ 3.002-3.002 $0.114$ $4.996$ $0.092$ 3.005-3.005 $0.163$
DGP B
$\lambda=0.7$ $0.451$ $0.034$ $0.700$ $0.015$ $0.701$ $0.022$
$\beta_{0,1}=1$ $0.990$ $0.074$ $1.000$ $0.047$ $1.000$ $0.048$
$\beta_{0,2}=1.5$ $1.441$ $0.136$ $1.000$ $0.047$ $1.500$ $0.084$
$\gamma_{0,1}=5$ $4.541$ $0.186$ $4.996$ $0.092$ $4.997$ $0.092$
$\gamma_{0,2}=-3$ 5.756-5.756 $0.512$ $4.996$ $0.092$ 3.005-3.005 $0.163$
DGP C
$\lambda=0.7$ $0.420$ $0.026$ $0.600$ $0.024$ $0.701$ $0.022$
$\beta_{0,1}=1$ $0.988$ $0.069$ $0.995$ $0.052$ $1.000$ $0.048$
$\beta_{0,2}=1.5$ $1.446$ $0.127$ $0.995$ $0.052$ $1.500$ $0.084$
$\gamma_{0,1}=5$ $4.426$ $0.192$ $4.769$ $0.108$ $4.997$ $0.092$
$\gamma_{0,2}=-3$ 5.423-5.423 $0.394$ $4.769$ $0.108$ 3.005-3.005 $0.163$
10% of Isolated Nodes
DGP A
$\lambda=0.7$ $0.699$ $0.013$ $0.699$ $0.013$ $0.699$ $0.020$
$\beta_{0,1}=1$ $1.002$ $0.045$ $1.001$ $0.045$ $1.001$ $0.046$
$\beta_{0,2}=1.5$ $1.503$ $0.084$ $1.504$ $0.084$ $1.503$ $0.085$
$\gamma_{0,1}=5$ $5.004$ $0.091$ $5.004$ $0.092$ $5.005$ $0.093$
$\gamma_{0,2}=-3$ 3.007-3.007 $0.099$ 3.006-3.006 $0.131$ 3.007-3.007 $0.165$
DGP B
$\lambda=0.7$ $0.465$ $0.032$ $0.699$ $0.013$ $0.699$ $0.020$
$\beta_{0,1}=1$ $0.983$ $0.082$ $1.001$ $0.045$ $1.001$ $0.046$
$\beta_{0,2}=1.5$ $1.425$ $0.154$ $1.504$ $0.084$ $1.503$ $0.085$
$\gamma_{0,1}=5$ $4.220$ $0.234$ $5.004$ $0.092$ $5.005$ $0.093$
$\gamma_{0,2}=-3$ 6.844-6.844 $0.490$ 3.006-3.006 $0.131$ 3.007-3.007 $0.165$
DGP C
$\lambda=0.7$ $0.421$ $0.021$ $0.567$ $0.023$ $0.699$ $0.020$
$\beta_{0,1}=1$ $0.980$ $0.075$ 50.032-50.032 $5.157$ $1.001$ $0.046$
$\beta_{0,2}=1.5$ $1.435$ $0.138$ $0.993$ $0.055$ $1.503$ $0.085$
$\gamma_{0,1}=5$ $4.047$ $0.235$ $1.496$ $0.096$ $5.005$ $0.093$
$\gamma_{0,2}=-3$ 6.230-6.230 $0.339$ $4.602$ $0.129$ 3.007-3.007 $0.165$
22% of Isolated Nodes (Add Health)
DGP A
$\lambda=0.7$ $0.700$ $0.013$ $0.700$ $0.013$ $0.699$ $0.020$
$\beta_{0,1}=1$ $1.000$ $0.046$ $1.000$ $0.046$ $1.000$ $0.046$
$\beta_{0,2}=1.5$ $1.502$ $0.086$ $1.502$ $0.086$ $1.502$ $0.086$
$\gamma_{0,1}=5$ $5.006$ $0.096$ $5.006$ $0.100$ $5.004$ $0.102$
$\gamma_{0,2}=-3$ 3.004-3.004 $0.090$ 3.003-3.003 $0.127$ 2.998-2.998 $0.183$
DGP B
$\lambda=0.7$ $0.480$ $0.034$ $0.700$ $0.013$ $0.699$ $0.020$
$\beta_{0,1}=1$ $0.975$ $0.097$ $1.000$ $0.046$ $1.000$ $0.046$
$\beta_{0,2}=1.5$ $1.419$ $0.172$ $1.502$ $0.086$ $1.502$ $0.086$
$\gamma_{0,1}=5$ $3.699$ $0.313$ $5.006$ $0.100$ $5.004$ $0.102$
$\gamma_{0,2}=-3$ 7.757-7.757 $0.518$ 3.003-3.003 $0.127$ 2.998-2.998 $0.183$
DGP C
$\lambda=0.7$ $0.415$ $0.021$ $0.525$ $0.024$ $0.699$ $0.020$
$\beta_{0,1}=1$ $0.970$ $0.081$ 43.385-43.385 $5.077$ $1.000$ $0.046$
$\beta_{0,2}=1.5$ $1.435$ $0.153$ $0.986$ $0.058$ $1.502$ $0.086$
$\gamma_{0,1}=5$ $3.442$ $0.307$ $1.492$ $0.110$ $5.004$ $0.102$
$\gamma_{0,2}=-3$ 6.718-6.718 $0.330$ $4.291$ $0.166$ 2.998-2.998 $0.183$
  • Model 1 accounts for a single fixed effect per school. Model 2 includes a single fixed effect per school with a dummy variable capturing isolated students. Model 3 is our structural model.
4  Empirical Illustration
In this section, we provide an empirical illustration of our econometric approach to estimating peer effects using a unique and now widely-used dataset from the National Longitudinal Study of Adolescent to Adult Health (Add Health). Specifically, our main objective in this section is to compare the estimate of peer effects obtained using our approach (equation 4) to that obtained using the classical linear-in-means peer effects specification (equation 5), in a context where isolated students are present.
4.1 Data
We use the Wave I in-school Add Health data, collected between September 1994 and April 1995. This is a dataset of a nationally representative sample of 90,118 students (7th to 12th grade) from 145 middle, junior high, and high schools across the US. It includes information on the social and demographic characteristics of students as well as their friendship links–in particular, their best friends, up to 5 females and up to 5 males.
After removing observations with missing data, the sample used for our empirical analysis encompasses 68,430 students from 141 schools. The number of students per school ranges widely from 18 to 2,027, with an average of 485 per school. On average, each student reports having 3.4 friends (1.6 male friends and 1.9 female friends). Moreover, there are 14,900 (22%) students who have no peers (isolated students), including 7,655 (11%) who are not fully isolated, that is, they are nominated by others. Only 1% of students nominate 10 friends, suggesting that the top coding issue resulting from students’ inability to nominate more than 10 friends is not a serious concern here.
The dependent variable, GPA, is the average grade across four subjects: mathematics, science, English/language arts, and history or social science. GPA is calculated on the basis of students’ grades in these four subjects, which we recoded as follows: A = 4, B = 3, C = 2, and D = 1. In our analysis, we include controls for several potential factors that may influence GPA. These factors include sex, age, Hispanic ethnicity, race, living arrangements (whether the student lives with both parents), duration of attendance at the current school, participation in school clubs, mother’s education level, and mother’s profession. We also control for contextual variables associated with the student’s social network by including the average of friends’ control variables.
Table LABEL:tab:stat provides definitions of the variables included in the empirical analysis and presents summary statistics. The average GPA of students is 2.8, while the average GPA of their friends is 2.9. On average, students are approximately 15 years old and have attended their current school for 2.5 years. In terms of demographic composition, the sample includes 48.7% boys, 16.4% Hispanics, 64.7% whites, 16.8% Blacks, 7% Asians, and 9.5% from other racial backgrounds. Furthermore, 74.1% of the students live with both parents. Regarding maternal education, about 30.6% of students’ mothers have attained a high school (HS) education level, while 16.9% have education levels beyond HS, and 41.9% have education levels below HS. In terms of mothers’ occupations, 20.2% of the students’ mothers work in professional occupations (such as teachers, doctors, lawyers, and executives), 20% are homemakers or do not work, and 43.5% hold other jobs. [^14] Add Health provides a detailed list with more than 15 categories. We combine these occupations into four broader categories along with a missing indicator.
As an important feature of the proposed method is to account for whether students are isolated or not isolated, we also present summary statistics for the subsample of students who have no friends. The average GPA for this group is slightly lower than that of students who have friends. We also observe that students who have no peers are more likely to be males and are often from minority groups, such as Hispanics, Blacks, and Asians.